OpenStudy (jenniferjuice):

Divide 36x3 - 96x2 + 48x by 12x. 3x2 - 8x + 4 3x2 + 8x - 4 3x3 - 8x + 4 3x3 + 8x - 4

OpenStudy (kirbykirby):

Take each term one by one and divide it by 12x Take the coefficient infront of the x's and divide that by 12. Then to divide out the x's, use the exponent rule: \[ \large \frac{x^a}{x^b}=x^{a-b}\]

OpenStudy (kirbykirby):

And for the signs, if you are just dividing by 12x, then all the signs remain the same .

OpenStudy (jenniferjuice):

3x3 - 8x2+4x

OpenStudy (jenniferjuice):

@kirbykirby

OpenStudy (kirbykirby):

\[ \large \frac{x^3}{x}=x^{3-1}=x^2\]

OpenStudy (jenniferjuice):

so the answer is a? @kirbykirby

OpenStudy (kirbykirby):

perfect :)

OpenStudy (jenniferjuice):

could u help me with one more? :)

OpenStudy (jenniferjuice):

@kirbykirby

OpenStudy (kirbykirby):

sure

OpenStudy (kirbykirby):

Here you have to do long division

OpenStudy (jenniferjuice):

i got \[2x2−4x−3+ 13/4x-3\]

OpenStudy (kirbykirby):

almost :), should be -13/(4x-3) at the end

OpenStudy (jenniferjuice):

thaks :D